OpenStudy (anonymous):

Solve the system 2x – 3y = 7 and 8x – 12y = 5 by using graph paper or graphing technology. What is the solution to the system?

OpenStudy (p0sitr0n):

2x = 3y +7 x = (3y +7)/2 substitute 8 ((3y +7)/2) -12y = 5 4(3y +7) -12y =5 12y +28 -12y =5 28 = 5 are you sure you copied well? if yes, the are no solutions